300 mm 300 mm Determine the internal forces at Point J when a=0. A B 480 mm 240 mm D 780 N
Answer #1

ACCORDING TO THE DATA AND DIAGRAM ARE GIVEN IN THE QUESTION,

Given , 2-500 А 1 By MA=0 By B By (300) = - 720 (780) N By = -1872 N 780N {fx=0; = 0 Bict 780 + A 5 ( Efy=0 By + 12 (A) = 0 1Internal forces es in J: H 2283. 62 N v = 613. 2 N M = (1444.6) (0.48) 1872 (0:3) M = 131.81 Nm